# Distributed Inter-Strand Coupling Current Model for Finite Element Simulations of Rutherford Cables

### v1

@@ 8 Conclusion and outlook
In this work, we presented a Distributed Inter-Strand Coupling Current (DISCC) model, which is a FE model that describes the IS coupling current dynamics of superconducting Rutherford cables without a need for individual strand discretization.
The DISCC model involves three scalar scaling parameters defining homogenized material properties. The values of these parameters need to be tuned based on reference solutions of a given cable geometry and properties. The tuning procedure of the scaling parameters is straightforward and can be performed without accounting for the internal strand dynamics. The internal strand dynamics (hysteresis, eddy, IF, and ohmic effects) can then be included in the homogenized model using the ROHM and ROHF models, as well as a current sharing equation for ohmic effects close to and above the strand critical current.
The resulting homogenized cable model contains all loss and magnetization contributions, captures their interdependency, reproduces differential inductance, and offers an important reduction of the computational effort with respect to conventional cable models. It also provides accurate solutions in strongly non-uniform field environments, in which analytical methods may have a limited accuracy.
@@ Acknowledgment
Appendix A Keystone angle and non-uniformities
Appendix B Models for internal strand dynamics
B.1 ROHM model
To describe the strand magnetization and the associated loss without discretizing the strand internal structure in detail, we use the ROHM model [16]. It consists of a chain of $N$ hysteretic cells, as shown in Fig. 29(a). Each cell $k\in\{1,\dots,N\}$ defines a relation between the magnetic field $\boldsymbol{h}$ and a magnetic flux density fraction $\boldsymbol{b}_{k}$, with weight $\alpha_{k}$. The magnetic flux density $\boldsymbol{b}$ is obtained as a weighted sum of the $\boldsymbol{b}_{k}$:
@@ B.1 ROHM model
The irreversibility parameters $\kappa_{k}$ and $\chi_{k}$ are defined as decreasing functions of $\|\boldsymbol{b}\|$ to account for the field dependence of the critical current density in superconducting filaments, and decreasing functions of the strand current to account for the reduced cross-section available for magnetization currents in presence of transport current. Finally, they are also function of the temperature. In this paper, for simplicity, we neglect the dependence of the parameters on the strand current and we consider a constant and uniform temperature.
@@
The parameter values used for the ROHM model in Sections 6 and 7 are given in Table 2. These values are representative of a strand similar to the strand in [16] at a temperature of $1.9$ K.
@@ B.2 ROHF model
The parameter values used for the ROHF model in the reference cable model of Sections 6 and 7 are given in Table 3. These values are representative of a strand similar to the strand in [16] at a temperature of $1.9$ K. The parameter values for the homogenized cable model are identical, except for the irreversibility parameters $\xi_{k}$ which are rescaled as described in Section 6.2.
@@
B.3 Current sharing model

### v2

@@ 8 Conclusion and outlook
In this work, we presented the Distributed Inter-Strand Coupling Current (DISCC) model. It is a FE model that describes the IS coupling current dynamics of superconducting Rutherford cables without a need for individual strand discretization.
The DISCC model involves three scalar scaling parameters defining homogenized material properties. The values of these parameters need to be tuned based on reference solutions of a given cable geometry and properties. The tuning procedure of the scaling parameters is straightforward and can be performed without accounting for the internal strand dynamics. The internal strand dynamics (hysteresis, eddy, IF, and ohmic effects) can then be included in the homogenized model using the ROHM and ROHF models, as well as a current sharing law for ohmic effects close to and above the strand critical current.
The resulting homogenized cable model contains all loss and magnetization contributions, captures their interdependency, reproduces changes in cable inductance, and offers an important reduction of the computational effort with respect to conventional cable models, with a minimum speed-up of $20$, while also being considerably simpler to implement. It provides accurate solutions in strongly non-uniform field environments, in which analytical methods may have a limited accuracy. In particular, the detailed description of the cable transient response under fast excitations, and the decomposition into distinct loss components, helps for the design of novel quench protection methods.
@@ Acknowledgment
Appendix A Models for internal strand dynamics
A.1 ROHM model
To describe the strand magnetization and the associated loss without discretizing the strand internal structure in detail, we use the ROHM model [19]. It is inspired by energy-based models for ferromagnetic materials [57, 58, 59].
It consists of a chain of $N$ hysteretic cells, as shown in Fig. 30(a). Each cell $k\in\{1,\dots,N\}$ defines a relation between the magnetic field $\boldsymbol{h}$ and a magnetic flux density fraction $\boldsymbol{b}_{k}$, with weight $\alpha_{k}$. The magnetic flux density $\boldsymbol{b}$ is obtained as a weighted sum of the $\boldsymbol{b}_{k}$:
@@ A.1 ROHM model
The irreversibility parameters $\kappa_{k}$ and $\chi_{k}$ are not constant, but decreasing functions of $b=\|\boldsymbol{b}\|$ to account for the field dependence of the critical current density $j_{\text{c}}(b)$ in superconducting filaments. The parametrization $j_{\text{c}}(b)$ is however not used directly as is, but instead replaced by smoothed versions of it, denoted as $f_{\kappa}(b)$ and $f_{\chi}(b)$, in order to reproduce the magnetization curves at low fields more accurately, as described in [19]. Temperature dependence can be directly included using the $j_{\text{c}}(b,T)$ scaling. We consider a uniform and constant temperature in this paper.
Finally, parameters $\kappa_{k}$ and $\chi_{k}$ should be decreasing functions of the strand transport current, in order to account for the reduced cross-section available for magnetization currents in presence of transport current. In this paper, for simplicity, we neglected this effect as its implementation for the reference model is not trivial (it requires coupling between global and local quantities). Its implementation for the homogenized model is straightforward.
@@
The parameter values used for the ROHM model in Sections 6 and 7 are given in Table 2. These values are representative of a strand similar to the strand in [19] at a temperature of $1.9$ K and were obtained based on simulations of detailed strand geometries following the procedure described in [19].
@@ A.2 ROHF model
The parameter values used for the ROHF model in the reference cable model of Sections 6 and 7 are given in Table 3. These values are representative of a strand similar to the strand in [19] at a temperature of $1.9$ K and were obtained based on simulations of detailed strand geometries following the procedure described in [60]. The parameter values for the homogenized cable model are identical, except for the irreversibility parameters $\xi_{k}$ which are rescaled as described in Section 6.2.
@@
A.3 Current sharing model